Obtain the sum – of – products and product - of - sum canonical form for π‘₯1⨁(π‘₯2 ⋆ π‘₯3 β€² )


1
Expert's answer
2021-11-12T08:20:39-0500

Let us obtain the sum-of-products and product-of-sum canonical form for "𝑥_1\oplus(𝑥_2 \cdot 𝑥_3')." For this let us construct the trush table:


"\begin{array}{||c|c|c||c|c|c||}
\hline\hline
x_1 & x_2 & x_3 & x_3' & x_2\cdot x_3' &𝑥_1\oplus(𝑥_2 \cdot 𝑥_3')\\
\hline\hline
0 & 0 & 0 & 1 & 0 & 0\\
\hline
0 & 0 & 1 & 0 & 0 & 0\\
\hline
0 & 1 & 0 & 1 & 1 & 1\\
\hline
0 & 1 & 1 & 0 & 0 & 0\\
\hline
1 & 0 & 0 & 1 & 0 & 1\\
\hline
1 & 0 & 1 & 0 & 0 & 1\\
\hline
1 & 1 & 0 & 1 & 1 & 0\\
\hline
1 & 1 & 1 & 0 & 0 &1\\
\hline\hline
\end{array}"


It follows from the trush table that the sum-of-products canonical form is the following:


"x_1'x_2x_3'+x_1x_2'x_3'+x_1x_2'x_3+x_1x_2x_3;"


and the product-of-sum canonical form is the following:


"(x_1+x_2+x_3)(x_1+x_2+x_3')(x_1+x_2'+x_3')(x_1'+x_2'+x_3)."